Cryptology Problems

To be discussed Wednesday, March 6, 1996.

1. Do problems 3.2, 4.1, 4.2, 4.3, and 4.4 in the textbook.

2. Let n = 143 be a modulus for use in RSA. Choose a public encryption exponent e and a private decryption exponent d which can be used with this modulus.